OpenStudy (anonymous):

Help Please!!!! supposed to find y coordinate plugging in x value into y=cos pix . first value is .4 but not sure if I just multiply .4 by -1 or multiply cos pi x .4?

OpenStudy (vishweshshrimali5):

\[\large{y = \cos(\pi*x)}\] Right ?

OpenStudy (anonymous):

that's what im confused on ?... because the problem tells me that the point (.5,0) lies on the curve y=cos pix and the next x coordinate is .4 but I'm not understanding if I multiply by -1 ?... or what cause I dont' understand how .5,0 is a point in the same graph

OpenStudy (vishweshshrimali5):

Sorry: No don't multiply it by -1

OpenStudy (vishweshshrimali5):

See... cos(pi x) is not always equal to -1... Instead for integral values of x: \[\large{\cos(\pi x) = (-1)^x}\]

OpenStudy (vishweshshrimali5):

But since here x is not integer so it is better to just multiply x by pi and then find its cos to get the answer

OpenStudy (anonymous):

ok. so actually I just take the -1 and raise it to the x value?

OpenStudy (vishweshshrimali5):

Yeah if x is an integer.... if it is not an integer then a big NO

OpenStudy (vishweshshrimali5):

Here x is 0.4 which is NOT an integer so you CANNOT use (-1)^x...

OpenStudy (anonymous):

oh. so... I just have to find it graphically ?

OpenStudy (vishweshshrimali5):

No ... you can use this method: \[\large{\cos(\pi x) = \cos(\pi * 0.4) = \cos(0.4 \pi)}\] Now you can use calculator to find out cos(0.4pi)

OpenStudy (vishweshshrimali5):

You can also find it graphically if you want but it would be a tad difficult

OpenStudy (anonymous):

so that would give me .99975...

OpenStudy (vishweshshrimali5):

How ?

OpenStudy (vishweshshrimali5):

See: \[\large{\cos(0.4\pi) = \cos(0.4*180)}\] (in degrees) \[\large{= \cos(72)}\] = 0.309016

OpenStudy (anonymous):

ooooohh... ok. i'm completely forgot degrees.. whoops.:(

OpenStudy (vishweshshrimali5):

:) Its okay

OpenStudy (anonymous):

thank you vishweshrimali5 you have opened my mind. :) much thanks.

OpenStudy (vishweshshrimali5):

Your welcome :)